OpenStudy (anonymous):

Solve for x square root of 8 x plus 1 equals 5

OpenStudy (anonymous):

\[\sqrt{8x +1} = 5\]

OpenStudy (anonymous):

8x + 1 = 25 8x = 24 x = 3

OpenStudy (anonymous):

no it equals 5 not 25

OpenStudy (anonymous):

When you take a "Square" on both sides, you get 5^2, which is equal to 25.